7 men can do a piece of work in 12 days. How many men will be required to do double the work in 8 days.
Found 2 solutions by mbgupta@hotmail.com, ramkikk66:
Answer by mbgupta@hotmail.com(1)
Answer by ramkikk66(644)

7 men can do a piece of work in 12 days. How many men will be required to do double the work in 8 days.

Ans:
7 men can do the work in 12 days.
1 man can do the work in (7*12) = 84 days
1 man can do "double" the work - takes double the time - in (84*2) = 168 days

So to complete 168 person-days of work in 8 days, you need
(168/8) = 21 men.

:)